OpenStudy (anonymous):

Find two numbers whose difference is 150 and whose product is a minimum.

OpenStudy (goldphenoix):

50, -50.

OpenStudy (anonymous):

difference "150" not hundred

OpenStudy (goldphenoix):

75, -75.

OpenStudy (anonymous):

ohh yeah thanks

OpenStudy (goldphenoix):

I remember having this question before, and the answer was 50, and -50.

OpenStudy (goldphenoix):

"Let x be the larger number Let x - 100 be the smaller number The difference between the two is 100. The product would be x(x-100) = x² - 100x. If you take the derivative, you get 2x - 100. Set this to zero: 2x - 100 = 0 2x = 100 x = 50 Alternatively, if you aren't familiar with calculus, turn the equation into vertex form for a parabola. Do this by completing the square: x² - 100x = 0 Take the coefficient on the x-term (-100), divide it by 2 (-50), and square it (2500). Add this to both sides: x² - 100x + 2500 = 2500 (x - 50)² = 2500 (x - 50)² - 2500 = 0 This is vertex form a(x - h)² - k = 0 and the vertex will be at the point x = h or x = 50. So either way you figure it, the numbers are 50 and -50. The product is -2500 which is minimum." - Puzzling